# Promoting the transition to quantum thinking: development of a secondary school course for addressing knowledge revision, organization, and epistemological challenges

### 31 Jan 2024 v6

@@ I Introduction
Another challenge faced both by introductory science students and physics majors enrolled in a QM course is the difficulty to overcome knowledge fragmentation, i.e., the organization of knowledge in small, disconnected pieces of contextual nature, which can be productively applied locally but which lack global coherence. This concerns, in the two cases, student interpretation of natural phenomena [5, 6], and their knowledge of the quantum model [7]. Research conducted at the end of upper-division QM courses and at the beginning of graduate instruction suggests that student reasoning is strongly context-dependent [8], and therefore that the development of a globally consistent knowledge structure may be only halfway even after prolonged periods of instruction. So far as we know, no investigation of this issue is available on secondary school students and non-physics-or-engineering majors who received traditional instruction on QM. However, the more limited scope of teaching/learning sequences (TLSs) designed for such student populations enhances the risk of promoting the construction of disconnected models valid only in the context of an individual phenomenon or experiment [9].
@@
Overall, a major goal of physics education research (PER) on QM is helping students overcome the manyfold challenges involved in learning QM. For this purpose, the PER community has produced in recent years a number of instructional materials and TLSs, drawing on different educational approaches. For instance, C. Singh and her coworkers built on their own research on common difficulties to design and revise interactive tutorials (QuILTs), with the aim to promote the construction of schemas consistent with QM principles (e.g., [13]). A recent review of educational research on QM at secondary and lower undergraduate level [14] divides the teaching approaches into those that focus on conceptual/mathematical understanding, on epistemic activities such as model-building, and on the interpretation of the theory. Favoring the development of an integrated conceptual understanding has been a basic aim of Malgieri et al., who implemented Feynman’s sum-over-paths approach by using GeoGebra simulations, so as to allow secondary school students to analyze different experimental setups with the same conceptual tools [9]. The TLS of Wittmann and Morgan for nonscience majors places special emphasis on personal epistemology (not to be confounded with scientific epistemology) as a means to help students work with nonintuitive content and to strengthen their understanding of scientific modelling [15]. Baily and Finkelstein made of the controversial nature of the interpretation of QM and of the discussion of students’ beliefs about it a topic onto itself, by designing a modern physics course for engineering majors aimed to help develop more consistent views of quantum phenomena, more sophisticated views of uncertainty, and greater interest in QM [16].
@@
Given the growing consensus in PER and science education at large to shift the focus from difficulties to student resources, i.e., pieces of prior knowledge that can be productively used in the learning process [17, 18], researchers are starting to ask how to put conceptual, symbolic and epistemological resources of students in the service of learning QM [19, 20]. However, as regards instructional materials on QM, there is a need to identify the possible links between specific sets of available knowledge elements or structures and possible educational strategies, and to empirically test their effectiveness.
@@
In order to find indications on how to address the aforementioned challenges and needs in the learning of QM, we chose to work in the framework of conceptual change. Since the release of The Structure of Scientific Revolutions by T. Kuhn [21], the theory change from CM to QM has been seen as an exemplary case of conceptual change in the history of science [22]. Educational research shows that this is a central element also behind the challenges students face in learning QM [23, 24, 14]. With the rise of educational models of conceptual change [12, 25], a number of researchers started to consider the problem of teaching QM as the design of strategies to effectively promote a conceptual change in individual learners [22, 27, 26, 9].
@@
In general, a conceptual change approach to science teaching can be described as “seeking to foster understanding and the adoption of scientific ideas as new systems of interpretation” [28] of natural phenomena. In the context of introductory science topics, challenges related to knowledge revision [29, 30, 5], fragmentation [5, 6], epistemology [28], as well as the use of resources [6] have been widely investigated in this framework. Indeed, educational models of conceptual change have been primarily developed to account for the transition from naïve to scientific knowledge, a process associated with the modification of conceptual structures formed in the context of lay culture. The transition from CM to QM, instead, requires changes in knowledge structures about a scientific theory, and developed as a result of instruction. While long-established models of conceptual change can represent a valuable aid also in addressing this transition, there is a need to develop approaches to the teaching of QM that take into account the differences involved in learning a successive theory.
@@
In addition, studies on conceptual change are undergoing a systemic turn, with a growing number of researchers coming to understand this process in terms of multiple interacting elements, which include, e.g., cognitive aspects, epistemology, social interaction [28]. Teaching-learning sequences on QM have prevalently focused on one or the other of these factors. Educational proposals based on a systemic approach, coordinating multiple factors at different levels of analysis, are currently lacking.
@@
In this paper, we describe the design and refinement of a QM course for secondary school based on a more systemic approach, which relies on an analysis of conceptual change in the learning of QM as a successive theory, and integrates cognitive and epistemological aspects including the examination and use of epistemic practices of theoretical physicists, and a careful approach to interpretive themes. Our goal is to address all the aforementioned challenges, taking advantage of available knowledge elements and intuition in the process. The analysis of conceptual change is aimed to provide an in-depth characterization of the first two challenges, suggesting how to leverage prior knowledge for achieving an effective revision, and how to identify conceptual tools suitable to make predictions on quantum processes across contexts. Student exploration of authentic practices in a theory-building activity aims to propose strategies for promoting the development of a plausible picture of the quantum world, and is assisted in this process by the development of an awareness of the foundational debate.
@@ II.1 Transition to quantum mechanics as a successive theory: cognitive aspects
    the exemplars of the theory: tasks and resolution strategies encountered in lectures, exercises, laboratory assignments, textbooks, etc. [33, p. 134].
Theory change is always accompanied by change in exemplars and in relevant concepts at different representational levels (new formation, evolution, disappearance [32]). Therefore, we need to consider not only ontological change in concepts, but also change in constructs used by the scientific community to represent these concepts, as well as the change in tasks and in resolution strategies. These features mark important differences with conceptual change processes at introductory level, since naïve science is neither socially shared nor mathematized.
@@ II.1.1 Change in concepts and constructs: addressing the challenge
This expression refers to physical quantities describing properties of systems and includes both dynamical variables, that in QM become observables, and parameters such as mass, that in non-relativistic QM behaves as a classical quantity. Its main pattern of evolution is denoted as categorical generalization (Fig. 1.a): for an aspect of the notion (attribute), some of the corresponding conceptual properties (values of the attribute) pertain to both CM and QM (gray boxes), the others only to QM (black boxes). In the light of this pattern, a strategy to promote a smooth revision may be the following: since the aspect involved includes at least a property that is valid both in CM and in QM, the instructor can start its discussion by fostering a productive use of classical intuition. Then, the new property(ies) can be made available to students by means of constructivist strategies (e.g., use of analogies for concepts, mathematical sense-making and/or embodied cognition for mathematical constructs). Finally, following the prevalence model of conceptual change [35], we rely on dissonance strategies to produce cognitive conflict, using the frame for structuring an end-of-unit table containing interpretive tasks to discriminate between the old and the new properties of the notion, identifying the correct context of application of each one .
The second example is the trajectory of the vector construct, that in CM is primarily used to represent physical quantities, while in QM it typically refers to the state of a system. Its conceptual dynamics presents only one pattern, denoted as value disjunction (Fig. 1.b): for an aspect of the notion, there are properties valid only in CM (white box) and properties valid only in QM (black box). As regards the aspect at hand, the quantum version of the notion is totally different from the classical one. The main distinction between the strategy suggested for categorical generalization and the strategy we suggest here is the need to start by building an understanding of the new property(ies), and to use prior intuition at the end, as a contrast [36].
@@
All this is summarized in our first design principle:
@@ II.1.2 Change in exemplars: addressing the challenge
The key is represented by decomposing a basic element of the inferential net of Fig. 2, i.e., the knowledge of the relation between the relevant observables, into binary relations between their values that arise in measurement. To this purpose, we borrow some language and elements from a specific perspective on the quantum theory, the Geneva-Brussels approach [see, e.g., 44]. In our course, a value of a system quantity - classical or quantum - that can be said to be either possessed by a physical system (when the probability to measure it is $1$) or not, is denoted as physical “property” and relations existing between values are denominated as “relations between properties”. The concept of property we use is a strongly restricted version of the original one, which includes not only values but also the union of disjoint intervals of values. Unless indicated otherwise, we will describe ideal measurements of discrete and continuous quantities only in terms of single values.
@@
This framework can be used as a conceptual basis for an educational reconstruction of quantum measurement across contexts and of the transition from the classical to the quantum measurement task. Since mutual unacquirability and incompatibility naturally arise in the exploration of spin or photon polarization measurements, these relations can be introduced in one of these contexts at an elementary level. Given the possibility to address measurement in a simple quantitative form (Malus’s law for photon polarization and its equivalent for spin), they can be justified to students as empirical regularities that are specific to quantum systems. Moving on to the relations between system quantities is almost immediate: two quantities are compatible if every property of each one is compatible with at least one property of the other, otherwise they are incompatible. Except in a limited number of cases (when two quantities are incompatible, but admit simultaneous eigenstates), relations between quantities can be qualitatively assessed in a similar way:
@@
Other contexts can be added to the picture either by using relations between properties or between observables. In both cases, the measurement process can be initially discussed at a qualitative level. If we opt for framing the issue in terms of relations between observables, by knowing which relations exist between the observables that initially have a definite value and the measured observable, we determine which of them are definite after the measurement and which not. In addition, the possession in advance or the lack of a property of the measured observable allows us to assess the nature of the process: respectively determinate or stochastic. This task can be accomplished also in the presence of degeneracy.
@@
As regards the transition from classical to quantum measurement, the relations represent a further instrument for addressing student’s need of comparability, since mutual unacquirability and compatibility can be expressed also in classical terms. In the classical regime, all system quantities are compatible with one another, and every point particle always possesses one property of each quantity. Thus, the emergence of incompatibility can be identified as an explanation of theory change with relation to measurement and the description of systems at a point in time.
As a result, we are able to formulate the second principle:
@@ II.2.1 Personal epistemology: theoretical modelling cycles
Personal epistemology may be introduced as an individual’s answers to questions such as “how do you know?” and “why do you believe?” [15]. Recent reviews on conceptual change and epistemic cognition report that there is a convincing body of research establishing a connection between more sophisticated epistemologies and deeper conceptual understanding in a particular domain [28, 45]. QM represents an ideal context for exploiting this synergy: a focus on epistemology may promote the learning of counterintuitive quantum content; on the other hand, a course of QM may be an opportunity for studying the practices of scientific modelling. Wittmann and Morgan, for instance, structured large part of their course around activities in which students work to build new concepts and create new knowledge, using lecture time to discuss and debate ideas in a peer-instruction format [15].
In order to put the aforementioned synergy in the service of learning QM, we too chose to focus on knowledge-building activities. However, given the wide range of possible activities of this kind, we endeavoured to identify the most appropriate ones for the content at hand. According to Sandoval et al., the conceptual, procedural, and epistemic expertise of a discipline is bound up in its specific practices [46]. But what practices characterize the construction of QM as a knowledge domain? A peek at the history of physics in the early 20th century suggests that theory-building is at the core of these practices. We concluded that involving students in theoretical modelling activities could be a promising strategy for helping them accept the quantum description of the world as a plausible and reliable product of their own inquiry, developing theoretical reasoning skills in the process. However, educational research on the epistemic practices that characterize the work of theoretical physicists is currently lacking. In Fig. 3, we propose a list of historically significant practices of theoretical nature used by physicists for building new scientific knowledge.
@@
The third principle underlying the design of our course is the following:
EPISTEMIC PRINCIPLE design the course around a modelling process that includes theoretical practices used by physicists in the historical development of the discipline, with the goal to help students accept the quantum description of the world as a plausible and reliable product of their own inquiry, thus promoting a smooth transition to a quantum perspective build theoretical reasoning skills
@@ II.2.2 Scientific epistemology: approach to interpretation
An additional interpretive choice concerns the wave-particle duality. Baily and Finkelstein adopt a “matter-wave perspective” [16], that allows students to interpret without paradoxes how a system can “know” whether two paths are open or only one of them in a “which-way” experiment. However, if the system propagates as a wave, students may ask what kind of medium supports or, equivalently, is perturbed by this wave. For this reason, in the construction of a full quantum model of a system, we adopt a field ontology, a perspective put forward in education also in recent years [e.g., 54].
@@ III Implementing the principles: development of the path and of the activities
In this section, we illustrate how the principles of design drove the development of the course and of individual activities within it. The content and the modelling process were informed by the interplay of the Principle of Knowledge Revision, the Principle of Knowledge Organization and the Epistemic Principle. The construction of the learning path is described in Section III.1. In the subsequent sections, each principle is addressed separately, showing how it guided the design of individual activities aimed to implement it. The first and the second principles are discussed respectively in Section III.2 and III.3. Converting the specific practices listed in Fig. 3 into authentic inquiry activities has been a particularly complex task that required the examination of different research perspectives and approaches (Section III.4.1). In Sections III.4.3 - III.4.5 we show how they were blended in the design of various types of activities. The impact of the Epistemological Principle on the design was conditioned by the chosen contexts and topics, and is addressed in Section III.5.

### 20 Feb 2024 v8

@@ I Introduction
Another challenge faced both by introductory science students and physics majors enrolled in a QM course is the difficulty to overcome knowledge fragmentation, i.e., the organization of knowledge in small, disconnected pieces of contextual nature, which can be productively applied locally but which lack global coherence. This concerns, in the two cases, student interpretation of natural phenomena [5, 6], and their knowledge of the quantum model [7]. Research conducted at the end of upper-division QM courses and at the beginning of graduate instruction suggests that student reasoning is strongly context-dependent [8], and therefore that the development of a globally consistent knowledge structure may be only halfway even after prolonged periods of instruction. So far as we know, no investigation of this issue is available on secondary school students and non-physics-or-engineering majors who received traditional instruction on QM. However, the more limited scope of teaching/learning sequences designed for such student populations enhances the risk of promoting the construction of disconnected models valid only in the context of an individual phenomenon or experiment [9].
@@
Overall, a major goal of physics education research on QM is helping students overcome the manyfold challenges involved in learning QM. A recent review of educational research on QM at secondary and lower undergraduate level [13] divides the teaching approaches into those that focus on conceptual/mathematical understanding, on epistemic activities such as model-building, and on the interpretation of the theory. Favoring the development of an integrated conceptual understanding has been a basic aim of Malgieri et al., who implemented Feynman’s sum-over-paths approach by using GeoGebra simulations, so as to allow secondary school students to analyze different experimental setups with the same conceptual tools [9]. The course of Wittmann and Morgan for nonscience majors places special emphasis on personal epistemology (not to be confounded with scientific epistemology) as a means to help students work with nonintuitive content and to strengthen their understanding of scientific modelling [14]. Baily and Finkelstein made of the controversial nature of the interpretation of QM and of the discussion of students’ beliefs about it a topic onto itself, by designing a modern physics course for engineering majors aimed to help develop more consistent views of quantum phenomena, more sophisticated views of uncertainty, and greater interest in QM [15].
@@
Given the growing consensus to shift the focus from difficulties to student resources, i.e., pieces of prior knowledge that can be productively used in the learning process [16, 17], researchers are starting to ask how to put conceptual, symbolic and epistemological resources of students in the service of learning QM [18, 19]. However, as regards instructional materials on QM, there is a need to identify the links between specific sets of available knowledge elements and possible educational strategies, and to empirically test their effectiveness.
@@
In order to find indications on how to address the aforementioned challenges and needs in the learning of QM, we chose to work in the framework of conceptual change (CC). Since the release of The Structure of Scientific Revolutions by T. Kuhn [20], the theory change from CM to QM has been seen as an exemplary case of CC in the history of science [21]. Educational research shows that this is a central element also behind the challenges students face in learning QM [22, 23, 13]. With the rise of educational models of CC [12, 24], a number of researchers started to consider the problem of teaching QM as the design of strategies to effectively promote a conceptual change in individual learners [21, 26, 25, 9].
@@
In general, a CC approach to science teaching can be described as “seeking to foster understanding and the adoption of scientific ideas as new systems of interpretation” [27] of natural phenomena. In the context of introductory science topics, challenges related to knowledge revision [28, 29, 5], fragmentation [5, 6], epistemology [27], as well as the use of resources [6] have been widely investigated in this framework. Indeed, educational models of CC have been primarily developed to account for the transition from naïve to scientific knowledge, a process associated with the modification of conceptual structures formed in the context of lay culture. The transition from CM to QM, instead, requires changes in knowledge structures about a scientific theory, and developed as a result of instruction. While long-established models of CC can represent a valuable aid also in addressing this transition, there is a need to develop approaches to the teaching of QM that take into account the differences involved in learning a successive theory.
@@
In addition, studies on CC are undergoing a systemic turn, with a growing number of researchers coming to understand this process in terms of multiple interacting elements, which include, e.g., cognitive aspects, epistemology, social interaction [27]. Teaching-learning sequences on QM have prevalently focused on one or the other of these factors. Educational proposals based on a systemic approach, coordinating multiple factors at different levels of analysis, are currently lacking.
@@
In this paper, we describe the design and refinement of a QM course for secondary school based on a more systemic approach, which relies on an analysis of CC in the learning of QM as a successive theory, and integrates cognitive and epistemological aspects. Our goal is to address all the aforementioned challenges, taking advantage of available knowledge elements and intuition in the process. The analysis of the cognitive aspects is aimed to provide an in-depth characterization of the first two challenges, suggesting how to leverage prior knowledge for achieving an effective revision, and how to identify conceptual tools suitable to make predictions on quantum processes across contexts. The epistemological analysis suggests to engage students in a theory-building activity, with the aim to propose strategies for promoting the development of a plausible picture of the quantum world, and is assisted in this process by the parallel development of an awareness of the foundational debate.
@@ II.1 Transition to quantum mechanics as a successive theory: cognitive aspects
    the exemplars of the theory: tasks and resolution strategies encountered in lectures, exercises, laboratory assignments, textbooks, etc. [32, p. 134].
Theory change is always accompanied by change in exemplars and in relevant concepts at different representational levels (new formation, evolution, disappearance [31]). Therefore, we need to consider not only ontological change in concepts, but also change in constructs used by the scientific community to represent these concepts, as well as the change in tasks and in resolution strategies. These features mark important differences with CC processes at introductory level, since naïve science is neither socially shared nor mathematized.
@@ II.1.1 Change in concepts and constructs: addressing the challenge
This expression refers to physical quantities describing properties of systems and includes both dynamical variables, that in QM become observables, and parameters such as mass, that in non-relativistic QM behaves as a classical quantity. Its main pattern of evolution is denoted as categorical generalization (Fig. 1.a): for an aspect of the notion (attribute), some of the corresponding conceptual properties (values of the attribute) pertain to both CM and QM (gray boxes), the others only to QM (black boxes). In the light of this pattern, a strategy to promote a smooth revision may be the following: since the aspect involved includes at least a property that is valid both in CM and in QM, the instructor can start its discussion by fostering a productive use of classical intuition. Then, the new property(ies) can be made available to students by means of constructivist strategies (e.g., use of analogies for concepts, mathematical sense-making and/or embodied cognition for mathematical constructs). Finally, following the prevalence model of CC [34], we rely on dissonance strategies to produce cognitive conflict. The frame is used for structuring an end-of-unit table containing interpretive tasks to discriminate between the old and the new properties of the notion, identifying the correct context of application of each one .
The second example is the trajectory of the vector construct, that in CM is primarily used to represent physical quantities, while in QM it typically refers to the state of a system. Its conceptual dynamics presents only one pattern, denoted as value disjunction (Fig. 1.b): for an aspect of the notion, there are properties valid only in CM (white box) and properties valid only in QM (black box). In this respect, the quantum version and the classical one are totally different. The main distinction between the strategy suggested for categorical generalization and the strategy we suggest here is the need to start by building an understanding of the new property(ies), and to use prior intuition at the end, as a contrast [35].
@@
All this is summarized in the first design principle:
@@ II.1.2 Change in exemplars: addressing the challenge
The key is represented by decomposing a basic element of the inferential net of Fig. 2, i.e., the knowledge of the relation between the relevant observables, into binary relations between their values, that arise in measurement. To this purpose, we borrow some language and elements from a specific perspective on the quantum theory, the Geneva-Brussels approach [see, e.g., 43]. In our course, a value of a system quantity - classical or quantum - that can be said to be either possessed by a physical system (when the probability to measure it is $1$) or not, is denoted as physical “property”, and relations existing between values are denominated as “relations between properties”. The concept of property we use is a strongly restricted version of the original one, which includes not only values but also the union of disjoint intervals of values. Unless indicated otherwise, we will describe ideal measurements of discrete and continuous quantities only in terms of single values.
@@
This framework can be used as a conceptual basis for an educational reconstruction of quantum measurement across contexts and of the transition from the classical to the quantum measurement task. Since mutual unacquirability and incompatibility naturally arise in the exploration of spin or photon polarization measurements, which can also be addressed in a simple quantitative form (Malus’s law for photon polarization and its equivalent for spin), these relations can be introduced and justified to students as empirical regularities that are specific to quantum systems. Moving on to the relations between system quantities is almost immediate: two quantities are compatible if every property of each one is compatible with at least one property of the other, otherwise they are incompatible. Except in a limited number of cases (when two quantities are incompatible, but admit simultaneous eigenstates), relations between quantities can be qualitatively assessed in a similar way:
@@
Other contexts can be added to the picture either by using relations between properties or between observables. If we opt for the latter, by knowing which relations exist between the observables that initially have a definite value and the measured observable, we determine which of them are definite after the measurement and which not. In addition, the possession in advance or the lack of a property of the measured observable allows us to assess the nature of the process: respectively determinate or stochastic. This task can be accomplished also in the presence of degeneracy.
@@
Finally, the relations represent an instrument for addressing student’s need of comparability with the classical measurement task. All of them except incompatibility can be identified also in the classical regime: all system quantities are compatible with one another; properties of the same quantity are mutually unacquirable; every point particle possesses one property of each quantity. Thus, the emergence of incompatibility represents an explanation of theory change with relation to the measurement task and the description of systems at a point in time.
We are now able to formulate the second principle:
@@ II.2.1 Personal epistemology: theoretical modelling cycles
Personal epistemology may be introduced as an individual’s answers to questions such as “how do you know?” and “why do you believe?” [14]. Recent reviews on CC and epistemic cognition report that there is a convincing body of research establishing a connection between more sophisticated epistemologies and deeper conceptual understanding in a particular domain [27, 44]. QM represents an ideal context for exploiting this synergy: a focus on epistemology may promote the learning of counterintuitive quantum content; conversely, a course of QM may be an opportunity for studying the practices of scientific modelling. Wittmann and Morgan, e.g., structured most of their course around activities in which students work to build new concepts and create new knowledge [14].
In order to put the aforementioned synergy in the service of learning QM, we chose to adopt a similar approach. However, given the wide range of possible knowledge-building activities, we endeavoured to identify the most appropriate ones for the content at hand. According to Sandoval et al., the conceptual, procedural, and epistemic expertise of a discipline is bound up in its specific practices [45]. But what practices characterize the construction of QM as a knowledge domain? A peek at the history of physics in the early 20th century suggests that theory-building is at the core of these practices. We concluded that involving students in theoretical modelling activities could be a promising strategy for helping them accept the quantum description of the world as a plausible product of their own inquiry, developing theoretical reasoning skills in the process. However, educational research on the epistemic practices that characterize the work of theoretical physicists is currently lacking. In Fig. 3, we propose a list of historically significant practices of this kind.
@@
EPISTEMIC PRINCIPLE design the course around a modelling process that includes theoretical practices used by physicists in the historical development of the discipline, with the goal to help students accept the quantum description of the world as a plausible product of their own inquiry, thus promoting a smooth transition to a quantum perspective build theoretical reasoning skills
@@ II.2.2 Scientific epistemology: approach to interpretation
An additional interpretive choice concerns the wave-particle duality. Baily and Finkelstein adopt a “matter-wave perspective” [15], that allows students to interpret without paradoxes how a system can “know” whether two paths are open or only one of them in a “which-way” experiment. However, if the system propagates as a wave, students may ask what kind of medium supports or, equivalently, is perturbed by this wave. For this reason, in the construction of a full quantum model of a system, we adopt a field ontology, a perspective put forward in education also in recent years [e.g., 53].
@@ III Implementing the principles: development of the path and of the activities
In this section, we illustrate how the principles of design drove the development of the course and of individual activities within it. The content and the modelling process were informed by the interplay of the Principle of Knowledge Revision, the Principle of Knowledge Organization and the Epistemic Principle. The construction of the learning path is described in Section III.1. In the subsequent sections, each principle is addressed separately, showing how it guided the design of individual activities aimed to implement it. The first and the second principles are discussed respectively in Section III.2 and III.3. Converting the specific practices listed in Fig. 3 into authentic inquiry activities has been a particularly complex task that required the examination of different research perspectives and approaches (Section III.4.1). In Sections III.4.2 - III.4.5 we show how they were blended in the design of various types of activities. The impact of the Epistemological Principle on the design was conditioned by the chosen contexts and topics, and is addressed in Section III.5.